# Polynomial extensions of Raimi’s theorem

### v2

@@ Abstract
Raimi’s theorem guarantees the existence of a partition of $\mathbb{N}$ into two parts with an unavoidable intersection property: for any finite coloring of $\mathbb{N}$, some color class intersects both parts infinitely many times, after an appropriate shift (translation). We establish a polynomial extension of this result, proving that such intersections persist under polynomial shifts in any dimension. Given non-constant polynomials $P^{(1)},\dots,P^{(f)}\in\mathbb{Z}[x]$ with positive leading coefficients and $P^{(j)}(0)=0$, we construct a partition of $\mathbb{N}^{k}$ such that for any coloring of $\mathbb{N}^{k}$ with finitely many colors, there exist $x_{0}\in\mathbb{N}$ and a single color class that meets all partition pieces after shifts by $x_{0}+P^{(j)}(h)$ in each of the $k$ coordinate directions, for every $j$ and infinitely many values $h\in\mathbb{N}$. Our proof exploits Weyl’s equidistribution theory, Pontryagin duality, and the structure of polynomial relation lattices. We also prove some finite analogues of the above results for abelian groups and $SL_{2}(\mathbb{F}_{q})$.
@@ 1 Introduction
There exists a subset $E\subseteq\mathbb{N}$ with the following property. For every partition of $\mathbb{N}$ into finitely many parts, $\mathbb{N}=\bigcup_{j=1}^{t}F_{j}$ with $t\in\mathbb{N},$ there exist $j\in\{1,2,\ldots,t\}$ and $k\in\mathbb{N}$ such that both $(F_{j}+k)\cap E$ and $(F_{j}+k)\setminus E$ are infinite sets.
In other words, after an appropriate shift $k$, the set $F_{j}+k$ unavoidably intersects both $E$ and its complement in infinitely many elements. This theorem was originally proved by Raimi using topological methods, and subsequently Hindman [13] provided an elementary proof. Strengthened versions specifying the densities of the partition sets or guaranteeing positive densities in the conclusion can be found in [1, 10].
@@
Let $r,t,k\in\mathbb{N}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $\mathbb{N}^{k}$ such that for any partition $\bigcup_{j=1}^{t}F_{j}$ of ${\mathbb{N}}^{k}$ into $t$ parts, there exist $m\in\{1,2,\ldots,t\}$, an element $x_{0}\in\mathbb{N}$, and a sequence $\{x_{n}\}_{n\geq 1}$ in $\mathbb{N}$ such that for every $h$ from the set $x_{0}+\mathrm{FS}(\{x_{n}\}_{n\geq 1})$ and every $i\in\{1,2,\ldots,r\}$, the set $(F_{m}+h\mathbf{1}_{k})\cap E_{i}$ has infinitely many elements.
@@
There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $\mathbb{N}^{k}$ such that for any partition $\bigcup_{j=1}^{t}F_{j}$ of ${\mathbb{N}}^{k}$, there exist $m\in\{1,\dots,t\}$, an element $x_{0}\in\mathbb{N}$, and a set $H\subset\mathbb{N}$ of positive lower density with the property that for every $h\in H$ and every $j\in\{1,\dots,f\}$, the set
@@
Let $r,t\in\mathbb{N}$. Let $G$ be a finite cyclic group of order $N$, written additively. There are constants $\alpha=\alpha(r,t)>0$ and $N_{0}=N_{0}(r,t)$ such that the following holds for every $N\geq N_{0}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $G$ such that for each partition $\bigcup_{j=1}^{t}F_{j}$ of $G$, there exist $m\in\{1,\dots,t\}$ and an element $h\in G$ such that
@@
Let $r,t\in\mathbb{N}$. There are constants $\alpha=\alpha(r,t)>0$ and $q_{0}=q_{0}(r,t)$ such that the following holds for every odd prime power $q>q_{0}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $SL_{2}(\mathbb{F}_{q})$ such that for every partition $\bigcup_{j=1}^{t}F_{j}$ of $SL_{2}(\mathbb{F}_{q})$, there exist $m\in\{1,\dots,t\}$ and an element $h\in SL_{2}(\mathbb{F}_{q})$ such that
@@ 1.1 Proof ideas and techniques
Notation. By $\{x_{n}\}_{n\geq 1}$ we mean a sequence with infinitely many elements $x_{1},x_{2},\ldots$, and if the sequence has $k$ elements, then we write $\{x_{n}\}_{n=1}^{k}$. For any $x\in\mathbb{R}$, let $\{x\}$ stand the fractional part of $x$. For $A\subseteq{\mathbb{N}}$, we define the lower density of $A$ as $\underline{d}(A)=\displaystyle\liminf_{N\to\infty}{|A\cap[1,N]|\over N}$, and the upper density of $A$ as $\overline{d}(A)=\displaystyle\limsup_{N\to\infty}{|A\cap[1,N]|\over N}$. If $\underline{d}(A)=\overline{d}(A)$, then we say the density of $A$ exists and denote it by $d(A):=\underline{d}(A)=\overline{d}(A)$.
@@ 2.1 Partition and an induced coloring
For $j\geq 0$ set
$S_{j}:=(1-2^{-j},\,1-2^{-(j+1)}]\subset(0,1),\quad a_{j}:=1-2^{-j},\quad w_{j}:=|S_{j}|=2^{-(j+1)}.$
@@
For $i\in\{1,\dots,r\}$ define the tile
@@
These $r$ tiles $T_{j,i}$ form a partition $S_{j}$.
Figure 1: The partition of $(0,1)$ with $M_{j}=1$ for all $j$
@@ 2.2 Preliminary lemmas
Let $\bigcup_{i=1}^{t}F_{i}$ be any partition of ${\mathbb{N}}^{k}$. There exist an $F_{m}$ and an interval $(x,y)\subseteq[0,1)$ such that $\varphi(F_{m})$ is everywhere dense in $(x,y)$.
@@
Indeed let’s assume that the opposite holds. Let $(x_{0},y_{0})\subseteq[0,1)$ any interval. Inductively let $j\in\{1,2,\ldots,t\}$. Assume that $\{\varphi(\mathbf{x})\colon\mathbf{x}\in F_{j}\}$ is not dense in $(x_{j-1},y_{j-1})$, hence there is an interval $(x_{j},y_{j})$ with $(x_{j},y_{j})\subseteq(x_{j-1},y_{j-1})$ and $\{\varphi(\mathbf{x})\colon x\in F_{j}\}\cap(x_{j},y_{j})=\emptyset$. After $t$ many steps, we get that $(x_{t},y_{t})\bigcap\bigcup_{j=1}^{t}\{\varphi(\mathbf{x})\colon\mathbf{x}\in F_{j}\}=\emptyset$, so $(x_{t},y_{t})\cap\{\varphi(\mathbf{x})\colon\mathbf{x}\in{\mathbb{N}}^{k}\}=\emptyset$. This leads to a contradiction. ∎
@@
Let $\mathbb{N}^{k}=\bigcup_{m=1}^{t}F_{m}$ be given. By Lemma 2.1, choose $m$ and an interval $J=(x,y)\subset(0,1)$ with $\varphi(F_{m})$ dense in $J$. Write $\delta:=y-x>0$.
@@
Then, $\bigcup_{i=1}^{r}\widetilde{E}_{i}$ forms a partition of $G$, and for every partition $\bigcup_{m=1}^{t}F_{m}$ of $G$, there exist an index $m^{\ast}\in\{1,\dots,t\}$ and a shift $h\in G$ such that
@@
Let $I_{1}\cup I_{2}\cup\cdot\cup I_{r}$ be a partition of $\mathbb{F}_{q}$ with $|I_{1}|=\cdots=|I_{r-1}|=\left\lfloor\frac{q}{r}\right\rfloor$ and $|I_{r}|=q-(r-1)\left\lfloor\frac{q}{r}\right\rfloor$.
@@
For each partition $\bigcup_{i=1}^{t}F_{i}$ of $SL_{2}(\mathbb{F}_{q})$, let $F_{m}$ be the set such that $|F_{m}|\geq q^{3}/2t$. Assume without loss of generality that the projection of $F_{m}$ onto the first two coordinates, denoted by $F_{m}^{1,2}$, is of size at least $q^{2}/4t$, and for each $(x,y)\in F_{m}^{1,2}$, the number of $z$ is at least $q/4t$.

### v3

@@ Abstract
Raimi’s theorem guarantees the existence of a partition of $\mathbb{N}$ into two parts with an unavoidable intersection property: for any finite coloring of $\mathbb{N}$, some color class intersects both parts infinitely many times, after an appropriate shift (translation). We establish a polynomial extension of this result, proving that such intersections persist under polynomial shifts in any dimension. Let $P^{(1)},\dots,P^{(f)}\in\mathbb{Z}[x]$ be non-constant polynomials with positive leading coefficients and $P^{(j)}(0)=0$ for every $j$. We construct a partition of $\mathbb{N}^{k}$ into an arbitrarily fixed finite number of pieces such that for any coloring of $\mathbb{N}^{k}$ with finitely many colors, there exist $x_{0}\in\mathbb{N}$ and a single color class that meets all partition pieces after shifts by $x_{0}+P^{(j)}(h)$ in each of the $k$ coordinate directions, for every $j$ and infinitely many values $h\in\mathbb{N}$. Our proof exploits Weyl’s equidistribution theory, Pontryagin duality, and the structure of polynomial relation lattices. We also prove some finite analogues of the above results for abelian groups and $SL_{2}(\mathbb{F}_{q})$.
@@ 1 Introduction
There exists a subset $E\subseteq\mathbb{N}$ with the following property. For every finite coloring of $\mathbb{N}$ with $t$ colors, $\mathbb{N}=\bigcup_{j=1}^{t}F_{j}$, $t\in\mathbb{N},$ there exist a single color class $F_{m}$, $1\leq m\leq t$, and $k\in\mathbb{N}$ such that both $(F_{m}+k)\cap E$ and $(F_{m}+k)\setminus E$ are infinite sets.
In other words, after an appropriate shift $k$, the set $F_{m}+k$ unavoidably intersects both $E$ and its complement in infinitely many elements. This theorem was originally proved by Raimi using topological methods, and subsequently Hindman [13] provided an elementary proof. Strengthened versions specifying the densities of the partition sets or guaranteeing positive densities in the conclusion can be found in [1, 10].
@@
Let $r,t,k\in\mathbb{N}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $\mathbb{N}^{k}$ such that for every finite coloring of $\mathbb{N}^{k}$ with $t$ colors, $\mathbb{N}^{k}=\bigcup_{j=1}^{t}F_{j},$ there exist a single color class $F_{m}$, $1\leq m\leq t$, an element $x_{0}\in\mathbb{N}$, and a sequence $\{x_{n}\}_{n\geq 1}$ in $\mathbb{N}$ such that for every $h$ from the set $x_{0}+\mathrm{FS}(\{x_{n}\}_{n\geq 1})$ and every $i\in\{1,2,\ldots,r\}$, the set $(F_{m}+h\mathbf{1}_{k})\cap E_{i}$ has infinitely many elements.
@@
There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $\mathbb{N}^{k}$ such that for every finite coloring of $\mathbb{N}^{k}$ with $t$ colors, $\mathbb{N}^{k}=\bigcup_{j=1}^{t}F_{j},$ there exist a single color class $F_{m}$, $1\leq m\leq t$, an element $x_{0}\in\mathbb{N}$, and a set $H\subset\mathbb{N}$ of positive lower density with the property that for every $h\in H$ and every $j\in\{1,\dots,f\}$, the set
@@
Let $r,t\in\mathbb{N}$. Let $G$ be a finite cyclic group of order $N$, written additively. There are constants $\alpha=\alpha(r,t)>0$ and $N_{0}=N_{0}(r,t)$ such that the following holds for every $N\geq N_{0}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $G$ such that for every finite coloring of $G$ with $t$ colors, $G=\bigcup_{j=1}^{t}F_{j},$ there exist a single color class $F_{m}$, $1\leq m\leq t$, and an element $h\in G$ such that
@@
Let $r,t\in\mathbb{N}$. There are constants $\alpha=\alpha(r,t)>0$ and $q_{0}=q_{0}(r,t)$ such that the following holds for every odd prime power $q>q_{0}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $SL_{2}(\mathbb{F}_{q})$ such that for every finite coloring of $SL_{2}(\mathbb{F}_{q})$ with $t$ colors, $SL_{2}(\mathbb{F}_{q})=\bigcup_{j=1}^{t}F_{j},$ there exist a single color class $F_{m}$, $1\leq m\leq t$, and an element $h\in SL_{2}(\mathbb{F}_{q})$ such that
@@ 1.1 Proof ideas and techniques
Notation. By $\{x_{n}\}_{n\geq 1}$ we mean a sequence with infinitely many elements $x_{1},x_{2},\ldots$, and if the sequence has $k$ elements, then we write $\{x_{n}\}_{n=1}^{k}$. For any $x\in\mathbb{R}$, let $\{x\}$ stand for the fractional part of $x$. For $A\subseteq{\mathbb{N}}$, we define the lower density of $A$ as $\underline{d}(A)=\displaystyle\liminf_{N\to\infty}{|A\cap[1,N]|\over N}$, and the upper density of $A$ as $\overline{d}(A)=\displaystyle\limsup_{N\to\infty}{|A\cap[1,N]|\over N}$. If $\underline{d}(A)=\overline{d}(A)$, then we say the density of $A$ exists and denote it by $d(A):=\underline{d}(A)=\overline{d}(A)$.
@@ 2.1 Partition and an induced coloring
For $j\geq 0$, set $S_{j}:=(1-2^{-j},\,1-2^{-(j+1)}]\subset(0,1),\quad a_{j}:=1-2^{-j},\quad w_{j}:=|S_{j}|=2^{-(j+1)}.$
@@
For $i\in\{1,\dots,r\}$, define the tile
@@
These $r$ tiles $T_{j,i}$ form a partition of $S_{j}$.
Figure 1: The partition structure of $(0,1)$ into bands $S_{j}$ and tiles $T_{j,i}$
@@ 2.2 Preliminary lemmas
Let $\bigcup_{i=1}^{t}F_{i}$ be any partition (coloring) of ${\mathbb{N}}^{k}$. There exist an $F_{m}$ and an interval $(x,y)\subseteq[0,1)$ such that $\varphi(F_{m})$ is everywhere dense in $(x,y)$.
@@
Assume that the opposite holds. Let $(x_{0},y_{0})\subseteq[0,1)$ be any interval. Inductively, let $j\in\{1,2,\ldots,t\}$. Assume that $\{\varphi(\mathbf{x})\colon\mathbf{x}\in F_{j}\}$ is not dense in $(x_{j-1},y_{j-1})$, hence there is an interval $(x_{j},y_{j})$ with $(x_{j},y_{j})\subseteq(x_{j-1},y_{j-1})$ and $\{\varphi(\mathbf{x})\colon x\in F_{j}\}\cap(x_{j},y_{j})=\emptyset$. After $t$ many steps, we get that $(x_{t},y_{t})\bigcap\bigcup_{j=1}^{t}\{\varphi(\mathbf{x})\colon\mathbf{x}\in F_{j}\}=\emptyset$, so $(x_{t},y_{t})\cap\{\varphi(\mathbf{x})\colon\mathbf{x}\in{\mathbb{N}}^{k}\}=\emptyset$. This leads to a contradiction. ∎
@@
For every finite coloring of $\mathbb{N}^{k}$ with $t$ colors, $\mathbb{N}^{k}=\bigcup_{m=1}^{t}F_{m}$. By Lemma 2.1, choose $m$ and an interval $J=(x,y)\subset(0,1)$ with $\varphi(F_{m})$ dense in $J$. Write $\delta:=y-x>0$.
@@
Then, $\bigcup_{i=1}^{r}\widetilde{E}_{i}$ forms a partition of $G$, and for every finite coloring of $G$ with $t$ colors, $G=\bigcup_{m=1}^{t}F_{m}$, there exist an index $m^{\ast}\in\{1,\dots,t\}$ and a shift $h\in G$ such that
@@
Let $I_{1}\cup I_{2}\cup\cdots\cup I_{r}$ be a partition of $\mathbb{F}_{q}$ with $|I_{1}|=\cdots=|I_{r-1}|=\left\lfloor\frac{q}{r}\right\rfloor$ and $|I_{r}|=q-(r-1)\left\lfloor\frac{q}{r}\right\rfloor$.
@@
For each finite coloring $\bigcup_{i=1}^{t}F_{i}$ of $SL_{2}(\mathbb{F}_{q})$ with $t$ colors, let $F_{m}$ be a set such that $|F_{m}|\geq q^{3}/2t$. Assume without loss of generality that the projection of $F_{m}$ onto the first two coordinates, denoted by $F_{m}^{1,2}$, is of size at least $q^{2}/4t$, and for each $(x,y)\in F_{m}^{1,2}$, the number of $z$ is at least $q/4t$.